OpenStudy (anonymous):

I WILL MEADEL AND FAN!!!!!!!!!!!!!!!!!!!!!!!! the distance between moutainview and crowntop is 6.5 inches on a map. if the scale on the map is 1inch=25miles, what is the actual distnce between the two towns. A.31.5 miles B.16.25 miles C.162.5 miles D. 165 miles

OpenStudy (anonymous):

If 1 inch = 25 miles, then 6.5 inches = x miles 1/25 = 6.5/x <----- Our proportion. Cross multiply. x = (25)(6.5) x = 162.5 miles

OpenStudy (anonymous):

If you want to give medals you need to learn how to spell medal correctly. Meadel? wtf?? really?

OpenStudy (anonymous):

THANKS AGIAN I GAVE U ANOTHER MEADEL KEEP HELPING ME AND YOULL HAVE 1BILLON MEADELS

OpenStudy (anonymous):

U NEED TO KNOW HOW TO PICK A NAME YOUR NAME IS GARBGE OpiGeode

OpenStudy (anonymous):

AND LEAVE ME ALONE IF U DID NOT WANT TO ANSWER MHY QUESTION YOU SHOULD NOT BE HERE ANY WAY

OpenStudy (anonymous):

MEDAL* NOT MEADAL GOD WTF

OpenStudy (anonymous):

Are you stupid?

OpenStudy (anonymous):

F!uCK U NAND LEAVE ME ALONE PLEASE